OpenStudy (anonymous):

solve (y /- 2) + 5 = −4 2 −2 18 −18

OpenStudy (kainui):

Where are you having problems with this? I can help you out.

OpenStudy (kainui):

Where do you get stuck

OpenStudy (dumbcow):

please use parenthesis if you write out the problem .... or use equation editor

OpenStudy (anonymous):

when you solve you get y-10=8 then y=18

OpenStudy (nurali):

(y /- 2) + 5 = −4 y/-2=-4-5 y/-2=-9 y=-2*-9 y=18